OpenStudy (anonymous):

Find all degree solutions. (Enter your answers in the form a + bk, where a is in [0, 180), b is the smallest possible positive number, and k represents any integer. Enter solutions from smallest to largest a - value.) 2\cos^{2} ({6}\theta) + 3\cos ({6}\theta)+1=0

OpenStudy (aum):

\[2\cos^{2} ({6}\theta) + 3\cos ({6}\theta)+1=0 \]

OpenStudy (aum):

This is a quadratic equation in cos^2. Factor and solve for theta.

OpenStudy (aum):

\[ 2\cos^{2} ({6}\theta) + 3\cos ({6}\theta)+1=0 \\ 2\cos^{2} ({6}\theta) + 2\cos ({6}\theta) + \cos ({6}\theta) +1=0 \\ 2\cos(6\theta)*(\cos(6\theta)+1) + 1(\cos(6\theta)+1) = 0 \\ (2\cos(6\theta)+1)*(\cos(6\theta)+1) = 0 \\ \cos(6\theta) = -\frac 12~~~\text{ or }~~~\cos(6\theta) = -1 \]Solve for theta.

OpenStudy (anonymous):

wow its u again!

OpenStudy (anonymous):

ur going to kill me. i put the wrong one up. i actually already resolved this one

OpenStudy (aum):

np. close the question.